OpenStudy (smurfy14):

Im so confused :( f(x)=(x^2+4)(3-2x)^2 Solve the inequality for a) f(x)>0 b) f(x)≥0 c) f(x)<0 d) f(x)≤0 Im so confused :( f(x)=(x^2+4)(3-2x)^2 Solve the inequality for a) f(x)>0 b) f(x)≥0 c) f(x)<0 d) f(x)≤0 @Mathematics

OpenStudy (anonymous):

ok first term is always positive because \[x^2\geq0\] and so \[x^2+4\geq 4\]

OpenStudy (anonymous):

so you can safely ignore that term completely

OpenStudy (anonymous):

\[(3-2x)^2>0\] for all x except \[x=\frac{3}{2}\] in which case you get 0

OpenStudy (smurfy14):

ignore what term and how do u know it is ≥ and not >?

OpenStudy (anonymous):

you can ignore \[x^2+4\] because it is always greater than or equal to 4, so it is always greater than 0

OpenStudy (smurfy14):

ok

OpenStudy (anonymous):

so the whole thing \[(x^2+4)(3-2x)^2\] is always greater than or equal to zero, and it is always greater then zero unless x = 3/2 now lets answer each question in turn.

OpenStudy (anonymous):

\[a) f(x)>0\] all x except 3/2

OpenStudy (smurfy14):

so x cannot equal 3/2 correct?

OpenStudy (anonymous):

for that one yes, because if x = 3/2 you get 0

OpenStudy (anonymous):

\[b) f(x)≥0 \] true for all x

OpenStudy (anonymous):

\[f(x)<0\] never true

OpenStudy (anonymous):

\[d) f(x)≤0 \] only true if x = 3/2

OpenStudy (smurfy14):

so d would be written as (-inf, 3/2]?